# Autonomous Robotic Bone Micro-Milling System with Automatic Calibration and 3D Surface Fitting

### v2

@@ Abstract
Automating bone micro-milling using a robotic system presents challenges due to the uncertainties in both the external and internal features of bone tissue. For example, during mouse cranial window creation, a circular path with a radius of 2 to 4 mm needs to be milled on the mouse skull using a microdrill. The uneven surface and non-uniform thickness of the mouse skull make it difficult to fully automate this process, requiring the system to possess advanced perceptual and adaptive capabilities. In this study, we address this challenge by integrating a Microscopic Stereo Camera System (MSCS) into the robotic bone micro-milling system and proposing a novel pre-measurement pipeline for the target surface. Starting from uncalibrated cameras, the pipeline enables automatic calibration and 3D surface fitting through a convolutional neural network (CNN)-based keypoint detection. Combined with the existing feedback-based system, we develop the world’s first autonomous robotic bone micro-milling system capable of rapidly, in real-time, and accurately perceiving and adapting to surface unevenness and non-uniform thickness, thereby enabling an end-to-end autonomous cranial window creation workflow without human assistance. Validation experiments on euthanized mice demonstrate that the improved system achieves a success rate of 85.7 % and an average milling time of 2.1 minutes, showing not only significant performance improvements over the previous system but also exceptional accuracy, speed, and stability compared to human operators.
Index Terms:
Robotics and Automation in Life Sciences, Computer Vision for Medical Robotics, Visual Servoing, Computer Vision for Automation.
@@ I Introduction
Bone micro-milling is an essential biological operation that involves using specialized milling tools to precisely remove or shape bone tissue at a microscale level [1]. This demands a high level of skill from the human operator due to the tiny scale and the need for high accuracy and safety. Automating this process with a robotic system could, in principle, make it more accurate and reproducible than by a human operator. However, this presents challenges for the perceptual and adaptive capabilities of the robotic system. The system needs to first accurately perceive the uncertainties in the characteristics of the bone tissue, including external features such as initial pose and surface flatness, as well as internal features like thickness and heterogeneity [2], and then generate milling trajectories that can adapt to these uncertainties.
For instance, in scientific experiments involving the transplantation of human organoids into mice and monitoring their growth [3], an important preparatory step is creating a circular cranial window with a radius of 2 to 4 mm on the mouse skull, whose thickness varies unevenly from 0.27 to 0.51 mm, depending on the strain, sex, and age of the mouse. To avoid damaging the fragile membrane beneath, skilled human operators typically use a handheld microdrill to mill the skull along a circular trajectory, ensuring all positions along the trajectory reach the same expected remaining thickness simultaneously under the observation of a high magnification microscope, and then remove the central bone flap with tweezers. However, automating this process using robotic system is very complex due to the surface of the mouse skull is uneven and its thickness varies. Additionally, these features dynamically change due to the potential skull deformation caused by contact with the drill bit. Thus, the system must not only possess autonomous and accurate perception capabilities for the uneven surface and non-uniform thickness of the mouse skull, but also be able to rapidly and in real time adapt to their dynamic changes during the milling process.
@@
Figure 1: The hardware setup of the autonomous robotic bone micro-milling system, including a robotic arm holding a microdrill, a head holder for mouse to be milled, and a MSCS.
Many studies developed robotic systems for automated cranial window creation and attempted to pre-scan the skull to obtain its surface information and thickness for milling trajectory planning. Pak et al. [4] obtains surface position and thickness by contact scanning, which involves drilling at pre-set test points while detecting changes in conductivity between the drill tip and the skull. Ghanbari et al. [5] used micro-CT, while Navabi et al. [6] combined optical coherence tomography (OCT) and machine learning, to obtain the skull anatomy. Jeong et al. [7] utilized an ultrashort pulsed Ti-sapphire oscillator to generate and detect second harmonic light for the metrology of the skull. Nonetheless, the above methods lack a feedback mechanism for real-time adaptation to the dynamic changes in the skull’s characteristics, and performing pre-scanning for each specimen may lead to a long overall duration.
@@
This study [4] [5] [6] [7] [8]
Direct surface measurement ×
Real-time feedback × × × ×
Short overall duration × × ×
Without human assistance × × ×
In the prior research by the present authors [8], an autonomous robotic bone micro-milling system was developed, which employed a convolutional neural network (CNN) to process input 2D images as real-time feedback to support adaptive updates of the milling trajectory. The performance of the system was evaluated by 20 eggshell milling experiments, achieving a success ratio of 80 % and an average milling time of 16.8 minutes. However, due to the lack of direct surface measurements, the initial milling trajectory can only be set to descend from the same height at all positions, resulting in inconsistent contact times with the surface across different locations. Since no feedback signal is available before contact occurs, the system is unable to adjust the trajectory effectively during this phase, leading to increased idle time, prolonged overall milling duration, and a higher likelihood of milling failures. Moreover, before autonomous milling begins, the drill bit must be manually teleoperated to a starting position where the distance between the bit and the surface is appropriate, requiring assistance from a skilled human operator.
In this study, we introduce a Microscopic Stereo Camera System (MSCS) (see Fig. 1) and propose a pre-measurement pipeline for the target surface through automatic calibration and 3D surface fitting. The pipeline does not require a time-consuming individual pre-scanning process, allowing rapid generation of an initial trajectory without human assistance. By integrating it into the previous system, the existing real-time feedback capability was retained while adding a direct measurement to enable early perception of uneven surfaces, representing the world’s first autonomous robotic bone micro-milling system that combines pre-measurement with a real-time feedback mechanism and achieving an end-to-end autonomous cranial window creation workflow, as shown in Fig. 2. The improvement effectively reduces the idle time before contact and ensures more consistent contact timing across different positions on the milling trajectory, thereby improving both efficiency and success rate of milling process. The contrast of the capabilities of this study with existing works is shown in Table I. The main contributions of this work can be summarized as follows:
@@
    A pre-measurement pipeline was proposed and validated that, with the introduction of MSCS, starts from uncalibrated cameras to reach automatic calibration and 3D surface fitting.
@@
    The pipeline was integrated with the system in the prior research, achieving an end-to-end automation workflow.
@@
    The effectiveness and performance of the entire improved system were validated through 7 trials of autonomous cranial window creation experiments.
II Hardware Setup
Figure 2: End-to-end automation workflow of the improved autonomous robotic bone micro-milling system. The parts in black (Completion Level Recognition and Trajectory Planner) are modified from the original system and will be introduced in Section IV, while the parts in red (Automatic Calibration and 3D Surface Fitting) are newly added in this work, which are the the main contribution and will be introduced in Section III. Additionally, the robot coordinate system {R} and the microscopic 3D coordinate system {Mi} are illustrated at the bottom of the figure.
The hardware setup of the system is shown in Fig. 1, utilizing a single-arm setup from the Multiarm Robotic Platform for scientific exploration [9]. The setup includes a robotic arm (CVR038, Densowave, Japan), a linear actuator and a circular rail actuator, which can be regarded as an 8-degree-of-freedom serial manipulator with joint values $q\in\mathbb{R}^{8}$. The robotic arm is equipped with a micro drill (MD1200, Braintree Scientific, USA) designed to mill the mouse skull secured in a head holder (SG-4N, NARISHIGE, Japan). To accommodate the confined workspace of the robotic system room, a modified MSCS setup was developed based on a mobile MSCS setup [10], consisting of two cameras (STC-HD853HDMI, Omron-Sentech, Japan) fitted with low-distortion macro lenses (VS-LDA75, VS Technology, Japan). The MSCS provides a reconstruction space of approximately $20\times 30\times 10\ \mathrm{mm}$ with a depth resolution of $0.10~{}\pm~{}0.02\ \mathrm{mm}$ that is sufficient for reconstructing the mouse skull [11], and a working distance of around $500\ \mathrm{mm}$ that meets the requirements for robotic movement [9].
III Proposed Pre-measurement Pipeline
The proposed pipeline aims to automatically pre-measure the target surface and generate a 3D-fitted initial trajectory $\boldsymbol{p}_{ini,i}$ that adapts to the surface unevenness for the initialization of milling trajectory. $\boldsymbol{p}_{ini,i}$ is a $n$-dimensional vector, with each element defined by
@@ III Proposed Pre-measurement Pipeline
where $i=1,2,...,n$, and $n$ refers to the the number of trajectory sampling points. The pipeline contains two parts, the Automatic Calibration (Section III-A) that determines the transformation from the microscopic 3D coordinate system {Mi} to the robot coordinate system {R} (see Fig. 2), and the 3D Surface Fitting (Section III-B) that samples trajectory points on the target surface at appropriate locations and transforms them to {R}.
Figure 3: The process of the Automatic Calibration.
III-A Automatic Calibration
With the introduction of MSCS, the point clouds of the target surface can be reconstructed with respect to {Mi} using the model in the prior research by the present authors [11]. Since the milling trajectory must be planned with respect to {R}, it is essential to determine the transformation matrix from {Mi} to {R} through calibration. However, due to the limited workspace at the microscale and insufficient features on the microdrill, the way of accessing the 3D positions of end-effector with respect to {Mi} in traditional hand-eye calibration methods become impractical, such as attaching Aruco markers [12] or checkerboard patterns [13], or identifying landmarks based on the geometric features [14]. Therefore, in this section, we propose a 2D keypoint detection-based Automatic Calibration, achieving an automatic, rapid, and real-time calibration process for microdrill without human assistance, as shown in Fig. 3. The process consists of 2D Keypoint Detection to detect the drill tip 2D positions using a CNN based on the input stereo RGB images, 3D Reconstruction to obtain 3D reconstructed drill tip points, and Landmark Calibration to solve the calibration problem between {Mi} and {R} by combining the calibration setpoints in {R} with the corresponding 3D reconstructed drill tip points in {Mi}.
III-A1 2D Keypoint Detection
To reconstruct the 3D points of the drill tip, its 2D projection coordinates in both the left and right images are required. This is primarily due to its small size, as directly accessing the reconstructed point clouds would lead to a suboptimal result. Therefore, a keypoint-detection CNN is utilized to obtain the drill tip 2D positions in both images. The CNN used is a minor adjustment of U-Net [15], where the adjustments include changing the input ground truths from segmentation masks to confidence maps of drill tip localization. By this, the drill tip 2D positions $\boldsymbol{p}_{l}\triangleq\begin{bmatrix}x_{l}&y_{l}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{2}$ (pixel) in the left image and $\boldsymbol{p}_{r}\triangleq\begin{bmatrix}x_{r}&y_{r}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{2}$ (pixel) in the right image can be obtained by locating the local maximum from the predicted confidence map.
III-A2 3D Reconstruction
@@ III-A Automatic Calibration
where $\boldsymbol{M}_{d}\in\mathbb{R}^{3\times 4}$ is the coefficient matrix and $\boldsymbol{b}_{d}\in\mathbb{R}^{3\times 1}$ is the bias, whose values can be calculated using the streamlined orthogonal projection model of disparity-depth mapping function (detailed in [11]):
@@
where $c_{x},c_{y}$ (pixel) are the camera’s principal point image coordinates, $P_{\rho x},P_{\rho y}$ (pixel/mm) respectively indicate the object’s pixel sizes in the $x$ and $y$ directions, $d_{e}$ (mm) is the effective working distance of the stereo microscope setup, and $h_{\rho}$ (pixel/mm) denotes the depth response to unit disparity in reality. More details on the calibration of the camera intrinsic parameters in (3) are provided in the supplementary document. By substituting (3) into (2), the 3D position of the reconstructed drill tip can be calculated.
III-A3 Landmark Calibration
@@
To obtain the corresponding landmarks, the controller sends a sequence of calibration setpoints to move the drill tip, represented as $\boldsymbol{p}_{i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,k\in\mathbb{N}$ with respect to {R}. For each setpoint, two images are captured from left and right cameras in MSCS, and its 3D position with respect to {Mi} can be reconstructed by (2). Thus, a sequence of 3D reconstructed drill tip points with respect to {Mi} can be obtained, represented as $\boldsymbol{p}_{i}^{\{\text{Mi}\}}\in\mathbb{R}^{3},i=1,\cdots,k\in\mathbb{N}$. Note that at least $k\geq 3$ non-collinear points are required to uniquely determine the transformation.
Finally, given the source landmarks $\boldsymbol{p}_{i}^{\{\text{Mi}\}}$ and the target landmarks $\boldsymbol{p}_{i}^{\{\text{R}\}}$, the rigid transformation $\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}$ in (4) can be calculated using the VTK’s classic Landmark Transform Method [^1] https://vtk.org/doc/nightly/html/classvtkLandmarkTransform.html, thereby completing the process of Automatic Calibration. The accuracy of calibration is then evaluated by the point-to-point distances after transformation (see Section V-C).
III-B 3D Surface Fitting
In this section, the goal is to generate an initial trajectory that is 3D fitted to the surface of the target, with its center point determined through the Trajectory Center Localization and trajectory points distributed on the circular path being sampled and biased through 3D-fitted Initial Trajectory Generation.
III-B1 Trajectory Center Localization
Figure 4: Schematic of bregma and lambda, which is expected to be detected through the CNN. The milling center can be calculated by linearly weighting the two points.
In the actual experiment of mice cranial window creation, the center of the circular milling trajectory is always located between two specific anatomical points on the skull: bregma and lambda, as illustrated in Fig. 4. Inspired by the method in [16], which uses a CNN to automatically detect bregma and lambda, we applied a U-Net-based CNN with the same architecture as the one in Section III-A1 for the detection of these two points.

### v3

@@ Abstract
Automating bone micro-milling using a robotic system presents challenges due to the uncertainties in both the external and internal features of bone tissue. For example, during mouse cranial window creation, a circular path with a radius of 2 to 4 mm needs to be milled on the mouse skull using a microdrill. The uneven surface and non-uniform thickness of the mouse skull make it difficult to fully automate this process, requiring the system to possess advanced perceptual and adaptive capabilities. In this study, we address this challenge by integrating a Microscopic Stereo Camera System (MSCS) into the robotic bone micro-milling system and proposing a novel online pre-measurement pipeline for the target surface. Starting from uncalibrated cameras, the pipeline enables automatic calibration and 3D surface fitting through a convolutional neural network (CNN)-based keypoint detection. Combined with the existing feedback-based system, we develop the world’s first autonomous robotic bone micro-milling system capable of rapidly, in real-time perceiving and adapting to surface unevenness and non-uniform thickness, thereby enabling an end-to-end autonomous cranial window creation workflow without human assistance. Validation experiments on euthanized mice demonstrate that the improved system achieves a success rate of 85.7 % and an average milling time of 2.1 minutes, showing not only significant performance improvements over the previous system but also exceptional accuracy, speed, and stability compared to human operators.
@@ I Introduction
Bone micro-milling is an essential biological operation that involves using specialized milling tools to precisely remove or shape bone tissue at a microscale level [10]. This demands a high level of skill from the human operator due to the tiny scale and the need for high accuracy and safety. Automating this process with a robotic system could, in principle, make it more accurate and reproducible than human operation. However, this presents challenges for the perceptual and adaptive capabilities of the robotic system. The system must first accurately perceive the uncertainties in the characteristics of bone tissue, including external features such as surface flatness, as well as internal features like thickness and heterogeneity [17], and generate milling trajectories that adapt to these uncertainties.
@@
Figure 1: The hardware setup of the autonomous robotic bone micro-milling system, including a robotic arm holding a microdrill, a head holder for the mouse to be milled, and a Microscopic Stereo Camera System (MSCS).
For instance, in scientific experiments involving the transplantation of human organoids into mice and monitoring their growth [18], an important preparatory step is creating a 2–4 mm circular cranial window on the mouse skull, whose thickness varies from 0.27 to 0.51 mm depending on strain, sex, and age. To avoid damaging the fragile underlying membrane, skilled operators use a handheld microdrill under a high-magnification microscope to mill along a circular trajectory, ensuring uniform remaining thickness, and then remove the central bone flap with tweezers. This procedure demands high operator skill, making automation essential for improving repeatability and robustness.
Recent studies have made significant progress in semi- or fully-automated craniotomy in medium-sized mammals. Xu et al. [19] developed a robotic-assisted system integrating human–robot collaboration and safe motion constraints, validated on Beagle models. Bian et al. [2] proposed a feed-force-monitoring algorithm for autonomous cranium milling, tested on in vitro dog and goat specimens. Zhichao et al. [22] introduced a computed tomography (CT)-based path planning method using the Virtual-Center approach, validated on living dogs. Compared with these animals, the mouse skull is significantly thinner and softer, with higher sensitivity to surface unevenness and non-uniform thickness. These characteristics impose stricter requirements on robotic precision, stability and safety, and potential skull deformation during milling introduces dynamic challenges for fully autonomous operation.
To address these challenges in mice, several approaches have attempted to employ offline measurement, which involves acquiring the skull anatomical information prior to milling through preoperative imaging or manual surface scanning. Pak et al. [15] used contact scanning, Ghanbari et al. [5] applied micro-CT, Navabi et al. [13] combined optical coherence tomography (OCT) with machine learning, and Jeong et al. [6] utilized ultrashort pulsed Ti-sapphire lasers for skull metrology. However, these offline measurement–based methods provide only static anatomical information of the skull, lack real-time feedback to adapt to dynamic changes, and require time-consuming measurements for each specimen, inevitably prolonging the overall operation time.
Other studies have attempted to achieve autonomous cranial window creation in mice using intraoperative feedback for real-time adaptation. Jia et al. [7] employed Bayesian transfer learning from manual eggshell milling to eggs of different sizes using images and force as feedback. In our previous work [21], we used CNN-based image feedback, achieving an 80% success rate and a milling time of 16.8 minutes in eggshell milling. Due to ethical constraints, these systems were validated only on eggshells, whose thickness ($\approx 0.4\ \mathrm{mm}$) and underlying delicate membrane are similar to those of the mouse skull [1]. However, in these feedback-based methods, the lack of pre-measured skull anatomical information prevents the initial trajectory from being pre-fitted to the target surface, requiring manual adjustment via teleoperation, which increases idle time, prolongs the overall milling duration, and raises the risk of failure. Moreover, the limited duration of inhalation anesthesia in mice (20–30 minutes [14]), the more complex structural characteristics of the mouse skull compared to eggshells, and the fragile underlying brain tissue make it difficult to directly apply these methods to live mice, where higher milling speed and success rate are required.
@@
This study [19] [2] [22] [15] [5] [13] [6] [7] [21]
Pre-measurement1 × ×
Real-time feedback × × × × ×
Short overall duration × × × × × ×
Without human assistance2 × × × × × × ×
Validation on mouse × × × × ×
  • 1
    Pre-measurement refers to obtaining skull anatomical information via offline measurement in [19, 2, 22, 15, 5, 13, 6], or via online measurement in this study.
  • 2
    Human assistance including scanning process in [19, 2, 22, 15, 13], or initial trajectory adjustment in [7, 21].
Figure 2: End-to-end automation workflow of the improved autonomous robotic bone micro-milling system: the red part, Proposed Online Pre-measurement Pipeline, including Automatic Calibration (Section II-A) and 3D Surface Fitting (Section II-B), is the main contribution of this work and responsible for surface measurement before milling; the black part, Intraoperative Real-time Feedback, including Completion Level Recognition (Section III-A) and Trajectory Planner (Section III-B), is improved from the original system [21] and responsible for closed-loop control during milling. “Robot” refers to the robot motion controller, which controls the robot’s end effector (microdrill) to reach the specified coordinates. The robot coordinate system {R} and the microscopic 3D coordinate system {Mi} are illustrated at the bottom.
More importantly, in the aforementioned methods, manual assistance is required, either during the scanning process in offline measurement-based systems or during the initial trajectory adjustment in feedback-based systems. This reliance on the operator’s skill compromises the system’s autonomy, efficiency, and reproducibility, and therefore these systems cannot be regarded as truly autonomous.
In this study, we develop a robotic system capable of perceiving the uneven surface and heterogeneous thickness of the mouse skull and adapting in real time, enabling fully autonomous cranial window creation. To this end, we introduce a Microscopic Stereo Camera System (MSCS, see Fig. 1) and propose an online pre-measurement pipeline. Unlike conventional offline measurement methods, which require a time-consuming scanning process and can only provide static anatomical information, this pipeline enables real-time reconstruction of the skull and automatically performs calibration and surface fitting, allowing rapid generation of an initial trajectory. By integrating this pipeline with the previous image-based intraoperative real-time feedback system, the improved automation workflow (see Fig. 2) is expected to enhance the overall performance and eliminate the need for human assistance. Based on these advances, we present the world’s first autonomous robotic bone micro-milling system, which combines pre-measurement with real-time feedback, enabling rapid and precise end-to-end creation of mouse cranial windows. Experiments on euthanized mice demonstrate that, compared to the previous system, our approach substantially increases both speed and success rate, demonstrating its potential applicability in future in vivo studies. The comparison of this study’s capabilities with existing works is presented in Table I. The main contributions of this work can be summarized as follows:
@@
    An online pre-measurement pipeline was proposed, which, with the introduction of MSCS, starts from uncalibrated cameras to achieve automatic calibration and 3D surface fitting.
@@
    An autonomous milling system was developed that synergizes the proposed pre-measurement pipeline with the image-based feedback in the prior research, achieving a rapid, precise, end-to-end automation workflow.
@@
    Validation experiments on euthanized mice were conducted, demonstrating excellent performance and efficiency even compared with skilled human operators.
II Proposed Online Pre-measurement Pipeline
The proposed pipeline aims to rapidly and automatically perform online pre-measurement of the target surface and generate a 3D-fitted initial trajectory $\boldsymbol{p}_{ini}$ that adapts to the surface flatness for the initialization of the milling trajectory. $\boldsymbol{p}_{ini}$ is a $n$-dimensional vector, with each element defined by
@@ II Proposed Online Pre-measurement Pipeline
where i=1,2,,ni=1,2,...,n, and $n$ refers to the number of trajectory sampling points. The pipeline contains two parts, the Automatic Calibration (Section II-A) that determines the transformation from the microscopic 3D coordinate system {Mi} to the robot coordinate system {R} (see Fig. 2), and the 3D Surface Fitting (Section II-B) that samples trajectory points on the target surface at appropriate locations and transforms them to {R}.
II-A Automatic Calibration
Figure 3: The Automatic Calibration determines the rigid transformation matrix $\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}$ from the microscopic 3D coordinate system {Mi} to the robot coordinate system {R}. The robot executes a sequence of predefined movement commands to reach calibration setpoints, at each of which the 3D reconstructed drill tip point is obtained through 2D Keypoint Detection and 3D Reconstruction. Landmark Calibration then uses these 3D points and setpoints to calculate $\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}$.
With the introduction of MSCS, the point clouds of the target surface can be reconstructed with respect to {Mi} using the model from the authors’ previous work [9]. Since the milling trajectory must be planned relative to {R}, it is necessary to calibrate the transformation matrix from {Mi} to {R}. However, at the microscale, the limited workspace and lack of features on the microdrill make traditional hand-eye calibration methods infeasible, such as using Aruco markers [4], checkerboard patterns [20], or geometric landmarks [11]. Therefore, in this section, we propose a keypoint detection-based Automatic Calibration, enabling rapid, real-time, and fully automated calibration, as shown in Fig. 3. The process includes 2D Keypoint Detection to detect 2D drill tip positions via a CNN, 3D Reconstruction to obtain 3D reconstructed drill tip points, and Landmark Calibration to solve the transformation between {Mi} and {R}.
II-A1 2D Keypoint Detection
To reconstruct the 3D points of the drill tip, its 2D projection coordinates in both the left and right images are required. This is primarily due to its small size, as directly accessing the reconstructed point clouds would lead to a suboptimal result. Therefore, a keypoint-detection CNN is utilized to obtain the drill tip 2D positions in both images. The CNN used is a minor adjustment of U-Net [16], where the adjustments include changing the input ground truths from segmentation masks to confidence maps of drill tip localization. By this, the drill tip 2D positions $\boldsymbol{p}_{l}\triangleq\begin{bmatrix}x_{l}&y_{l}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{2}$ in the left image and $\boldsymbol{p}_{r}\triangleq\begin{bmatrix}x_{r}&y_{r}\end{bmatrix}^{\text{T}}\in\mathbb{R}^{2}$ in the right image can be obtained by locating the local maximum from the predicted confidence map.
II-A2 3D Reconstruction
@@ II-A Automatic Calibration
where $\boldsymbol{M}_{d}\in\mathbb{R}^{3\times 4}$ is the coefficient matrix and $\boldsymbol{b}_{d}\in\mathbb{R}^{3\times 1}$ is the bias, whose values can be calculated using the streamlined orthogonal projection model of disparity-depth mapping function following the methodology presented in [9]:
@@
where $c_{x},c_{y}$ (pixel) are the camera’s principal point image coordinates, $P_{\rho x},P_{\rho y}$ (pixel/mm) respectively indicate the object’s pixel sizes in the $x$ and $y$ directions, $d_{e}$ (mm) is the effective working distance of the stereo microscope setup, and $h_{\rho}$ (pixel/mm) denotes the depth response to unit disparity in reality. More details on the 3D reconstruction model and its calibration are provided in the Supplementary Materials. By substituting (3) into (2), the 3D position of the reconstructed drill tip can be calculated.
II-A3 Landmark Calibration
@@
To obtain the corresponding landmarks, $k$ calibration setpoints are sent to the robot to move the microdrill tip to these positions, whose coordinates with respect to {R} are represented as $\boldsymbol{p}_{i}^{\{\text{R}\}}\in\mathbb{R}^{3},i=1,\cdots,k$. For each setpoint, two images are captured from left and right cameras in MSCS, and its 3D position with respect to {Mi} can be reconstructed by (2). Thus, $k$ reconstructed drill tip points with respect to {Mi} can be obtained, represented as $\boldsymbol{p}_{i}^{\{\text{Mi}\}}\in\mathbb{R}^{3},i=1,\cdots,k$. Note that at least 3 non-collinear points ($k\geq 3$) are required to uniquely determine the transformation.
Finally, given the source landmarks $\boldsymbol{p}_{i}^{\{\text{Mi}\}}$ and the target landmarks $\boldsymbol{p}_{i}^{\{\text{R}\}}$, the transformation $\boldsymbol{T}_{\text{\{Mi\}}\rightarrow\text{\{R\}}}$ in (4) can be calculated using the VTK’s classic Landmark Transform Method [^1] https://vtk.org/doc/nightly/html/classvtkLandmarkTransform.html, thereby completing the Automatic Calibration.
II-B 3D Surface Fitting
In this section, the goal is to generate an initial trajectory that is 3D-fitted to the target surface. The center point of the trajectory is determined through the Trajectory Center Localization, while the trajectory points are distributed along the circular path and further offset through the 3D-fitted Initial Trajectory Generation. A schematic is illustrated in Fig. 4.
II-B1 Trajectory Center Localization
Figure 4: Schematic of 3D Surface Fitting. The bregma and lambda are detected through a CNN, and the milling center $\boldsymbol{p}_{c}^{\{\text{Mi}\}}$ is calculated by linearly weighting these two points. Subsequently, surface-fitted trajectory $\boldsymbol{p}_{fit}^{\{\text{Mi}\}}$ are sampled on the target surface based on $\boldsymbol{p}_{c}^{\{\text{Mi}\}}$ and radius $r$, and the 3D-fitted initial trajectory $\boldsymbol{p}_{ini}^{\{\text{Mi}\}}$ is obtained by offsetting $\boldsymbol{p}_{fit}^{\{\text{Mi}\}}$ along the normal vectors $\boldsymbol{n}$ by a distance $r_{d}+\Delta d$.
In the actual experiment of mouse cranial window creation, the center of the circular milling trajectory is always located between two specific anatomical points on the skull: bregma and lambda. Inspired by the method in [23], which uses a CNN to automatically detect bregma and lambda, we applied a U-Net-based CNN with the same architecture as in Section II-A1 for the detection of these two points.